OpenStudy (anonymous):

If you invest $7700 for six years, are you better off with an interest rate of 4% compounded quarterly or 3.7% compounded continuously?

OpenStudy (anonymous):

the answer is quarterly but i got a higher number for continuously.

OpenStudy (anonymous):

can someone work this out so i can see what i did wrong?

OpenStudy (anonymous):

here are the formulas: p(1+r/n)^nt (is the regular one) pe^rt (is the continuous one)

OpenStudy (turingtest):

P=7700 t=6 quarterly: n=4 r=0.04 \[P(1+r/n)^{nt}=7700(1+0.04/4)^{4*6}\approx9777\] continuous: r=0.037 \[Pe^{rt}=7700e^{0.036*6}\approx9556\] so the answer should be quarterly. I don;t know where you went wrong.

OpenStudy (anonymous):

I used the same rate for both like a dummy

OpenStudy (anonymous):

Thanks!